OpenStudy (anonymous):

A farmer plans to fence a rectengular pasture adjacent to a river. The pasture must contain 245,000 square meters to provide enough grass for the herd, What dimensions will require the least amount of fencing if no fencing is needed along the river. (Calc, Optimization)

OpenStudy (anonymous):

L * w = 245,000 m² -> L = 245,000 / w Fence: 2 w + L f = 2w + 245,000 / w Let f' = 0 to obtain minimum amount of fence needed: @Fey Can you continue?

OpenStudy (anonymous):

I need a few minutes to process this x)

OpenStudy (anonymous):

Okie, as long as I know you're following :)

OpenStudy (anonymous):

Why is L 245,000 and what does w mean?

OpenStudy (anonymous):

oh! Nevermind got it!